40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 21 19:28:03 2021, Anchal, Update, PSL, PSL/IFO recovery 
    Reply  Wed Apr 21 19:43:20 2021, Koji, Update, PSL, New HEPA speed control P_20210421_193637.jpgP_20210421_193627.jpg
       Reply  Thu Apr 22 14:41:55 2021, Chub, Update, PSL, New HEPA speed control 40M_PSL_HEPA.jpg
          Reply  Fri Apr 23 18:00:02 2021, gautam, Update, PSL, HEPA speed lowered HEPAdiag.pdf
             Reply  Fri Apr 23 19:26:58 2021, Koji, Update, PSL, HEPA speed lowered 
             Reply  Fri May 14 17:45:05 2021, rana, Update, PSL, HEPA speed raised 
                Reply  Tue May 18 00:52:38 2021, rana, Update, PSL, HEPA speed raised Untitled.png
                   Reply  Tue May 18 20:26:11 2021, rana, Update, PSL, HEPA speed raised 
    Reply  Thu Apr 22 15:15:26 2021, gautam, Update, PSL, PMC transmission 
       Reply  Thu Apr 22 15:34:54 2021, Anchal, Update, PSL, PMC transmission 
Message ID: 16076     Entry time: Thu Apr 22 15:15:26 2021     In reply to: 16068     Reply to this: 16077
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC transmission 

I was a bit surprised by these numbers suggesting the PMC transmission is only 50-60%. I went to the table today and confirmed that it is more like 85% (1.3 W in, 1.1 W transmitted, both numbers from with the FieldMate power meter), as I claimed in 2019. Even being conservative with the power meter errors, I think we can be confident T_PMC will be >80% (modulo any thermal effects with higher power degrading the MM). There isn't any reliable record of what the specs of the PMC mirrors are, but assuming the IO couplers have T=4000ppm and the end mirror has T=500ppm as per Alan's plot, this is consistent with a loss of something like 300ppm loss per mirror - seems very high given the small beam spots, but maybe these mirrors just aren't as high quality as the test masses?

It's kind of unfortunate that we will lose ~20% of the amplifier output through the first filter, but I don't see an easy way to clean these mirrors. It's also not clear to me if there is anything to be gained by attempting a cleaning - isn't the inside of the cavity supposed to be completely isolated from the outside? Maybe some epoxy vaporization events degraded the loss?

Quote:

The transmitted power was ~50-60 mW. (Had to use power meter suspended by hand only.

ELOG V3.1.3-